# Integrating Large Language Models into Recommendation via Mutual Augmentation and Adaptive Aggregation

### v1

@@ Abstract.
Conventional recommendation methods have achieved notable advancements by harnessing collaborative or sequential information from user behavior. Recently, large language models (LLMs) have gained prominence for their capabilities in understanding and reasoning over textual semantics, and have found utility in various domains, including recommendation. Conventional recommendation methods and LLMs each have their own strengths and weaknesses. While conventional methods excel at mining collaborative information and modeling sequential behavior, they struggle with data sparsity and the long-tail problem. LLMs, on the other hand, are proficient at utilizing rich textual contexts but face challenges in mining collaborative or sequential information. Despite their individual successes, there is a significant gap in leveraging their combined potential to enhance recommendation performance.
@@
In this paper, we introduce a general and model-agnostic framework known as Large Language model with mutual augmentation and adaptive aggregation for Recommendation (Llama4Rec). Llama4Rec synergistically integrates conventional and LLM-based recommendation models. Llama4Rec proposes data augmentation and prompt augmentation strategies tailored to enhance the conventional model and the LLM respectively. An adaptive aggregation module is adopted to combine the predictions of both kinds of models to refine the final recommendation results. Empirical studies on three real-world datasets validate the superiority of Llama4Rec, demonstrating its consistent and significant improvements in recommendation performance over baseline methods.
${}^{\dagger}$oresponding Author
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14-18, 2024; Washington D.C., USA
1. Introduction
@@
Recommender systems have emerged as crucial solutions for mitigating the challenge of information overload (He et al., 2023a, b; Luo et al., 2023a, 2022a). Recommender systems encompass a multitude of tasks, such as rating prediction (Steck, 2013; Khan et al., 2021) and top-$k$ recommendation (Le and Lauw, 2021; Luo et al., 2023b). The top-$k$ recommendation, which encompasses collaborative filtering-based direct recommendation (He et al., 2017, 2020), sequential recommendation (Kang and McAuley, 2018; Sun et al., 2019; Chen et al., 2023), and more, has found wide applications in various areas. However, recommender systems still suffer from the data sparsity and long-tail problem. Data sparsity arises from sparse user-item interactions, making the task of accurately capturing user preferences more challenging. The long-tail problem further intensifies data sparsity issue, as a substantial number of less popular items (i.e., long-tail items) are infrequently interacted with, leading to inadequate data for effective model training and compromised recommendation quality.
@@
In recent years, Large Language Models (LLMs) have emerged, exhibiting exceptional capabilities in language understanding, text generation, and complex reasoning tasks (OpenAI, 2023; Touvron et al., 2023a, b; Wang et al., 2023; Zhou et al., 2023). Recent studies have started exploring their applicability in recommender systems (Liu et al., 2023; Bao et al., 2023; Zhang et al., 2023c). For example, Liu et al. employed ChatGPT with in-context learning (ICL) for various recommendation tasks (Liu et al., 2023). Further progress has been achieved by adopting the instruction tuning technique (Ouyang et al., 2022; Longpre et al., 2023) to align general-purpose LLMs with recommendation tasks for improved performance (Zhang et al., 2023c; Bao et al., 2023). For instance, TALLRec (Bao et al., 2023) reformulates the recommendation problem as a binary classification task and introduces an effective instruction fine-tuning framework for adapting the LLaMA model (Touvron et al., 2023b). However, these LLM-based recommendation methods may not perform optimally as they do not harness the collaborative or sequential information captured by conventional recommendation models.
@@
Conventional recommendation models and LLM-based recommendation methods each have their respective strengths and weaknesses. Conventional methods excel in mining collaborative information and modeling sequential behaviors, while LLMs are proficient in leveraging rich textual contexts. As such, the integration of LLMs into recommender systems presents a significant opportunity to amalgamate the advantages of both methodologies while circumventing their respective shortcomings. There have been initial attempts to harness the strengths of both conventional and LLM-based recommenders (Geng et al., 2022; Zhang et al., 2023b; Zheng et al., 2023; Wei et al., 2023). Some efforts have sought to integrate collaborative/sequential information by enabling LLMs to comprehend user/item ID information (Geng et al., 2022; Zhang et al., 2023b; Zheng et al., 2023). For instance, a concurrent study by Zhang et al. encoded the semantic embedding into the prompt (Zhang et al., 2023b) and send it to LLM. On the other hand, some research works have aimed to augment conventional models using LLMs via data or knowledge augmentation (Wei et al., 2023; Xi et al., 2023). LLMRec (Wei et al., 2023) enhances recommender systems by deploying LLMs to augment the interaction graph, thereby addressing the challenges of data sparsity and low-quality side information.
@@
However, existing methods have several limitations. Firstly, current methods lack generalizability. The strategy of integrating ID information proves challenging to generalize across different domains and necessitates additional training. The current data augmentation method is not universally applicable, as it only addresses a limited number of recommendation scenarios. Secondly, current research primarily focuses on the integration at the data-level (e.g., data augmentation) or model-level (e.g., make LLM understand ID semantics), leaving the result-level integration largely unexplored. Lastly, there is an absence of a comprehensive framework that combines and integrates these methods into a single construct. In light of these limitations, our objective is to explore the integration of conventional recommendation models and LLM-based recommendation methods in depth to address the above limitations and enhance recommendation performance.
@@
In this paper, we introduce a general framework known as Large language model with mutual augmentation and adaptive aggregation for Recommendation, referred to as Llama4Rec, for brevity. The core idea of Llama4Rec is to allow conventional recommendation models and LLM-based recommendation models to mutually augment each other, followed by an adaptive aggregation of the augmented models to yield more optimized results. Specifically, Llama4Rec performs data augmentation for conventional recommendation models by leveraging instruction-tuned LLM to alleviate the data sparsity and long-tail problem. The data augmentation is tailored with different strategies depending on the recommendation scenarios. Furthermore, we use conventional recommendation models to perform prompt augmentation for LLMs. The prompt augmentation includes enriching collaborative information from similar users and providing prior knowledge from the conventional recommendation model within the prompt. We also propose an adaptive aggregation module that merges the predictions of the LLM and conventional models in an adaptive manner. This module is designed as a simple yet effective way to combine the strengths of both models and refine the final recommendation results. We conduct empirical studies on three real-world datasets, encompassing three different recommendation tasks, to validate the superiority of our proposed method. The results consistently demonstrate its superior performance over baseline methods, highlighting notable improvements in recommendation performance.
@@
  • We introduce Llama4Rec, a general and model-agnostic framework to integrate LLM into conventional recommendation models. Llama4Rec performs the data augmentation for conventional models to alleviate the data sparsity problem and improve model performance. The prompt augmentation is applied to LLM for leveraging the information captured by the conventional models.
  • Llama4Rec employs an adaptive aggregation approach to combine the prediction from the conventional recommendation model and LLM for improved recommendation performance via leveraging and merging the information by both kinds of models.
@@
2. Related Work
2.1. Conventional Recommendation Methods
@@
Conventional recommendation methods serve as the cornerstone for the contemporary landscape of recommender systems (Zhang et al., 2019). Representative recommendation tasks include rating prediction, collaborative filtering-based direct recommendation, and sequential recommendation, where the latter two are usually formulated as top-$k$ recommendation problems. Specifically, one of the seminal techniques is the use of matrix factorization for rating prediction, popularized by methods such as Singular Value Decomposition (SVD) (Koren et al., 2009). Collaborative filtering (CF) is another commonly used technique for the recommender systems (Adomavicius and Tuzhilin, 2005). Recent advancements have evolved CF techniques into more complex neural network architectures and graph-based models (He et al., 2017; Wang et al., 2019) to enhance the model performance. Sequential recommendation models incorporate temporal patterns into the recommendation pipeline. Techniques such as recurrent neural networks have been adapted for this purpose (Hidasi et al., 2015). Recent research focuses on applying attention mechanisms to further refine these models, leading to a noteworthy boost in performance (Sun et al., 2019; Kang and McAuley, 2018).
@@
Although conventional recommendation techniques are well-suited for capturing latent information associated with users and items, they often require a substantial amount of user-item interactions to provide accurate recommendations, which limits their effectiveness in data sparse and long-tail scenarios (Park and Tuzhilin, 2008).
@@
Figure 1. (i) The overall framework architecture of the proposed Llama4Rec consists of two main components: mutual augmentation and adaptive aggregation. The mutual augmentation includes data augmentation and prompt augmentation. (ii) Illustration of the data augmentation process encompasses three diverse recommendation scenarios. (iii) The pipeline of the adaptive aggregation module, which merges the predictions from the conventional recommendation model and the LLM.
2.2. Large Language Model for Recommendation
LLMs have brought a paradigm shift in numerous areas of machine learning, including recommendation methods (Fan et al., 2023). One of the most compelling advantages of LLM-based recommendation methods is their capacity for contextual understanding and in context learning (Dong et al., 2022). Inspired by this, reference (Liu et al., 2023) utilized ChatGPT across diverse recommendation tasks and found it to be effective in specific contexts, underpinned by robust experiments and human evaluations. Similarly, Wang et al. (Wang and Lim, 2023) introduced a novel zero-shot technique for next-item recommendations, further substantiating the utility of LLMs in this arena. However, it is noteworthy that these methods do not consistently demonstrate a marked improvement over conventional recommendation algorithms, which is largely attributable to the inherent misalignment between the general-purpose capabilities of LLMs and the specialized requirements of the recommendation task. To address this issue, recent studies further attempt to instruct tuning the specific LLM to align with human preference (Bao et al., 2023; Zhang et al., 2023c). Typically, these approaches involve creating an instruction tuning dataset in line with recommendation tasks, which is then used to tune the LLM for recommendation. Such methodologies have demonstrated improved performance in generating more aligned and accurate recommendations.
Nevertheless, while LLMs excel at capturing intricate textual patterns, they may encounter challenges in comprehensively encoding user and item collaborative or sequential information. Though some concurrent studies (Zhang et al., 2023b; Zheng et al., 2023) aim to address this gap, they often lack in terms of generalizability and comprehensibility. In response to this challenge, we propose a novel framework designed to mitigate this issue.
3. Preliminary
We consider a recommender system with a set of users, denoted $\mathcal{U}=\{u_{1},u_{2},\ldots,u_{n}\}$, and a set of items, denoted =$\mathcal{I}=\{i_{1},i_{2},\ldots,i_{m}\}$. The rating prediction task aims to estimate the unknown values of $r_{ui}$ in the user-item interaction matrix $R\in\mathbb{R}^{n\times m}$, where each entry $r_{ui}$ is the rating assigned by user $u$ to item $i$. Different from rating prediction, top-$k$ recommendation focuses on identifying a subset of items $\mathcal{S}_{u}\subset\mathcal{I}$ for each user $u$. The subset is chosen to maximize a user-specific utility $U(u,\mathcal{S})$ with the constraint $|\mathcal{S}|=k$, which can be formally expressed as:
(1) $\mathcal{S}_{u}={\arg\max}_{\mathcal{S}\subset\mathcal{I},|\mathcal{S}|=k}U(u,\mathcal{S}).$
In the context of LLM-based recommendation methods, let ${LLM}$ represent the original LLM. These kinds of methods first utilize prompt to interpret the recommendation task into natural language. The LLM-based recommendation for user $u$ with in-context learning is denoted by ${r}_{u}={LLM}(\mathcal{P}_{u})$ where $\mathcal{P}_{u}$ is the recommendation prompt for user $u$. The recommendation prompt could either ask LLM to predict a rating towards a target item, or rank candidate items derived by retrieval models in top-k recommendation. To instruction fine-tune LLM, a dedicated dataset $\mathcal{D}_{ins}$ consisting of various instructions is utilized. The resulting instruction-tuned LLM is denoted as ${LLM}^{\prime}$. Therefore, the recommendation process in the fine-tuned model can be succinctly represented as $r_{u}={LLM}^{\prime}(\mathcal{P}_{u})$.
4. Methodology
4.1. Overview
@@
Figure 1 depicts the architecture of Llama4Rec, which consists of three components: data augmentation, prompt augmentation, and adaptive aggregation. More specifically, Llama4Rec leverages an instruction-tuned LLM to enhance conventional recommendation systems through data augmentation. The specific data augmentation strategies for different recommendation situations are detailed in Section 4.2. In addition, we employ conventional recommendation models to augment the LLM via prompt augmentation, with details in Section 4.3. To further refine the predictions of the conventional model and the LLM, we propose a simple yet effective adaptive aggregation module in Section 4.4. Lastly, we describe the training strategy for LLM in Section 4.5.
4.2. Data Augmentation for Conventional Recommendation Model
@@
We design ad-hoc data augmentation strategies for different recommendation scenarios to mitigate prevalent issues of data sparsity and the long-tail problem. This design is motivated by the fact that data distribution and tasks significantly vary across different recommendation scenarios. In the context of direct recommendation, we capitalize on the power of the instruction-tuned LLM to predict items that a user may like or dislike. We form pairs of these items to calculate the Bayesian Personalized Ranking (BPR) (Rendle et al., 2012) loss. For sequential recommendation, we harness the capabilities of the instruction-tuned LLM to predict items that are highly preferred by the user. These predicted items are then randomly inserted into the sequence of items the user has interacted with. For rating prediction, we utilize the LLM to extract valuable side information (i.e., missing attributes), which is then seamlessly integrated as additional features within the training data.
4.2.1. Data Augmentation for Direct Recommendation

### v2

@@ Abstract
Conventional recommender systems and Large Language Model (LLM)-based recommender systems each have their strengths and weaknesses. While conventional recommendation methods excel at mining collaborative information and modeling sequential behavior, they struggle with data sparsity and the long-tail problem. LLM, on the other hand, is proficient at utilizing rich textual contexts but faces challenges in mining collaborative or sequential information. Despite their individual successes, there is a significant gap in leveraging their ensemble potential to enhance recommendation performance.
@@
In this paper, we introduce a general and model-agnostic framework known as Large language models with mutual augmentation and adaptive aggregation for Recommendation (Llama4Rec), aiming to bridge this gap via explicitly ensemble LLM and conventional recommendation model for more effective recommendation. We propose data augmentation and prompt augmentation strategies tailored to enhance the conventional recommendation model and LLM respectively. An adaptive aggregation module is adopted to combine the predictions of both kinds of models to refine the final recommendation results. Empirical studies on three datasets validate the superiority of Llama4Rec, demonstrating significant improvements in recommendation performance.
I Introduction
@@
Recommender systems have emerged as crucial solutions for mitigating the challenge of information overload [68, 37, 35, 38, 39, 32]. Recommender systems encompass a multitude of tasks, such as rating prediction [48, 21, 36] and top-$k$ recommendation [24, 64, 34]. The top-$k$ recommendation, which encompasses collaborative filtering-based direct recommendation [15, 14], sequential recommendation [20, 49, 4], and more, has found wide applications in various areas. However, recommender systems still suffer from the data sparsity and long-tail problem. Data sparsity arises from sparse user-item interactions, making the task of accurately capturing user preferences more challenging. The long-tail problem further intensifies the data sparsity issue, as a substantial number of less popular items (i.e., long-tail items) are infrequently interacted with, leading to inadequate data for effective model training and compromised recommendation quality.
Figure 1: An example that illustrates the motivation of Llama4Rec.
@@
In recent years, Large Language Models (LLMs) have emerged, exhibiting exceptional capabilities in language understanding, text generation, and complex reasoning tasks [41, 51, 52, 2]. Recent studies have started exploring their applicability in recommender systems [29, 3, 66]. For example, Liu et al. employed ChatGPT with in-context learning (ICL) for various recommendation tasks [29]. Further progress has been achieved by adopting the instruction tuning technique [42, 31] to align general-purpose LLMs with recommendation tasks for improved performance [66, 3]. For instance, TALLRec [3] reformulates the recommendation problem as a binary classification task and introduces an effective instruction fine-tuning framework for adapting the LLaMA model [52]. However, these LLM-based recommendation methods may not perform optimally as they do not harness the collaborative or sequential information captured by conventional recommendation models.
Conventional recommendation models and LLM-based recommendation methods each have their respective strengths and weaknesses. An illustrative example is shown in Figure 1. Conventional methods excel in mining collaborative information and modeling sequential behaviors, while LLMs are proficient in leveraging rich textual contexts.
@@
Model ensemble has emerged as a powerful technique in the field of reliable machine learning [8, 75]. One of the key advantages of model ensemble is its ability to improve model performance. Each model in the ensemble may capture different aspects of the underlying data patterns, leading to a more comprehensive and robust problem understanding. By aggregating the predictions of these models, we can obtain a final prediction that is often more accurate and reliable than that of any single model. Inspired by the concept of model ensemble, the integration of LLMs into recommender systems presents a significant opportunity to amalgamate the advantages of both methods while circumventing their respective shortcomings. There have been initial attempts to harness the strengths of both conventional and LLM-based recommender systems [11, 69, 72, 57]. Some efforts have sought to integrate collaborative or sequential information by enabling LLMs to comprehend user/item ID information [11, 69, 72]. For instance, a concurrent study by Zhang et al. encodes the semantic embedding into the prompt and sends it to LLM [69]. Besides, Zheng et al. [72] proposes adapting large language models for recommendation by integrating collaborative semantics, combining contextual understanding with collaborative filtering to enhance recommendation performance. On the other hand, some research works have aimed to augment conventional models using LLMs via data or knowledge augmentation [57, 61]. For example, LLMRec [57] enhances recommender systems by deploying LLMs to augment the interaction graph, thereby addressing the challenges of data sparsity and low-quality side information.
@@
However, existing methods have several limitations. Firstly, current methods lack generalizability. Current strategy of integrating ID information in LLMs proves challenging to generalize across different domains and necessitates additional training. Moreover, the current data augmentation method is not universally applicable, as it only addresses a limited number of recommendation scenarios. Secondly, current research primarily focuses on the integration at the data-level (e.g., data augmentation) or model-level (e.g., making LLM understand ID semantics), leaving the result-level integration largely unexplored. Lastly, there is an absence of a comprehensive framework that combines and integrates these methods into a holistic solution. In light of these limitations, our objective is to explore the integration of conventional recommendation models and LLM-based recommendation methods in depth to address the above limitations and enhance recommendation performance.
@@
In this paper, we introduce a generic framework named Large language model with mutual augmentation and adaptive aggregation for Recommendation, referred to as Llama4Rec, for brevity. The core idea of Llama4Rec is to allow conventional recommendation models and LLM-based recommendation models to mutually augment each other, followed by an adaptive aggregation of the augmented models to merge the textual and collaborative modality and yield more optimized results. Specifically, Llama4Rec performs data augmentation for conventional recommendation models by leveraging instruction-tuned LLM to alleviate the data sparsity and long-tail problem and thus enrich the collaborative modality. The data augmentation is tailored with different strategies depending on the recommendation scenarios. Furthermore, we use conventional recommendation models to perform prompt augmentation for LLMs, enhancing the textual modality. The prompt augmentation includes enriching collaborative information from similar users and providing prior knowledge from the conventional recommendation model within the prompt. We also propose an adaptive aggregation module that merges the predictions of the LLM and conventional models in an adaptive manner. This module is designed as a simple yet effective way to combine the strengths of both models and refine the final recommendation results. We conduct empirical studies on three real-world datasets, encompassing three different recommendation tasks, to validate the superiority of our proposed method. The results consistently showcase its superior performance compared to baseline methods, achieving on average 14.65% improvement in direct recommendation, a 14.21% improvement in sequential recommendation, and a 3.10% improvement in rating prediction.
@@
  • We introduce Llama4Rec, a general and model-agnostic framework to integrate LLM into conventional recommendation models, merging the textual modality with collaborative modality. Llama4Rec performs the data augmentation for conventional models to alleviate the data sparsity problem and improve model performance. The prompt augmentation is applied to LLM for leveraging the information captured by the collaborative-modal recommendation models.
  • Llama4Rec employs an adaptive aggregation approach to combine the prediction from the conventional recommendation model and LLM for improved recommendation performance via leveraging and merging the information captured by such kinds of models working on different modalities.
@@
II Related Work
II-A Conventional Recommendation Methods
@@
Conventional recommendation methods serve as the cornerstone for the contemporary landscape of recommender systems [68]. Representative recommendation tasks include rating prediction, collaborative filtering-based direct recommendation, and sequential recommendation, where the latter two are usually formulated as top-$k$ recommendation problems. Specifically, one of the seminal techniques is the use of matrix factorization for rating prediction, popularized by methods such as Singular Value Decomposition (SVD) [22]. Collaborative filtering (CF) is another commonly used technique for the recommender systems [1]. Recent advancements have evolved CF techniques into more complex neural network architectures and graph-based models [15, 56] to enhance the model performance. On the other hand, sequential recommendation models incorporate temporal patterns into the recommendation pipeline. Techniques such as recurrent neural networks have been adapted for this purpose [16]. Recent research focuses on applying attention mechanisms to further refine these models, leading to a noteworthy boost in performance [49, 20].
@@
Although conventional recommendation techniques are well-suited for capturing latent information associated with users and items, they often require a substantial amount of user-item interactions to provide accurate recommendations, which limits their effectiveness in data sparse and long-tail scenarios [43].
@@
Figure 2: (i) The overall framework architecture of the proposed Llama4Rec consists of two main components: mutual augmentation and adaptive aggregation. The red dashed lines denoted the data augmentation and prompt augmentation process, respectively. (ii) Illustration of the data augmentation process, encompassing three diverse recommendation scenarios. (iii) The pipeline of the adaptive aggregation module, which includes the aggregation of predictions from both types of models. The symbols $u$, $i$, and $a$ represents user, item, and attributes, respectively.
II-B Large Language Model for Recommendation
Large Language Models (LLMs) have ushered in a paradigm shift across various domains, including recommender systems [70, 10]. Their ability to perform contextual understanding and reasoning offers significant advantages for recommendation tasks [27, 7]. Existing approaches leveraging LLMs in recommender systems can be broadly categorized into two groups: LLMs as recommenders and LLMs as enhancers.
The first category involves adapting LLMs to directly serve as recommenders, typically through prompting [5], in-context learning [54, 29], or fine-tuning [3, 66, 33]. For instance, Liu et al. [29] explored the use of ChatGPT across diverse recommendation tasks, demonstrating its effectiveness in specific contexts through rigorous experiments and human evaluations. Similarly, Bao et al. [3] proposed TALLRec, which fine-tunes LLMs using instruction tuning datasets tailored to recommendation tasks, aligning LLMs with recommendation objectives. The second category employs LLMs as enhancers to augment traditional recommender systems by generating or extracting semantic features, such as item descriptions or user profiles [58, 60, 53, 71, 45]. For example, Wei et al. [58] introduced LLMRec, which integrates LLMs with graph augmentation to enrich item representations for recommendation. By leveraging the semantic capabilities of LLMs, these approaches enhance the performance of conventional recommenders, offering a complementary strategy to improve recommendation quality.
In contrast to prior work, which typically adopts either the LLM as recommender or LLM as enhancer approach, our method uniquely combines the strengths of both paradigms. By integrating LLMs as both direct recommenders and enhancers within a unified framework, our approach leverages contextual understanding and semantic feature extraction to achieve superior recommendation performance.
TABLE I: Summary of notations used in this work.
Notation Definition
$u,i$ Users and items
$\hat{y}_{ui}$ Predicted score of user $u$ for item $i$
$e_{u},e_{i}$ Embeddings of user $u$ and item $i$
$\mathcal{D}$ Training dataset
$\mathcal{A}$ Set of attributes
$\mathcal{P}$ Prompt for the LLM
$\mathcal{L}_{\text{BPR}}$ Bayesian personalized ranking (BPR) loss
$\ell_{u}$ Long-tail coefficient for user $u$
$N(u)$ Number of interactions for user $u$
$U$ Utility score
$\Theta$ Original parameters of the LLM
$\alpha_{1},\alpha_{2}$ Hyperparameters
III Methodology
III-A Overview
@@
Figure 2 depicts the architecture of Llama4Rec, which consists of three components: data augmentation, prompt augmentation, and adaptive aggregation. More specifically, Llama4Rec leverages an instruction-tuned LLM to enhance conventional recommender systems through data augmentation. The specific data augmentation strategies for different recommendation situations are detailed in Section III-B. In addition, we employ conventional recommendation models to augment the LLM via prompt augmentation, with details in Section III-C. To further refine the predictions of the conventional model and the LLM, we propose a simple yet effective adaptive aggregation module in Section III-D. Lastly, we describe the training strategy for LLM in Section III-E. Summary of notations are listed in Table I.
III-B Data Augmentation for Conventional Recommender System
@@
We design ad-hoc data augmentation strategies for different recommendation scenarios to mitigate prevalent issues of data sparsity and the long-tail problem. This design is motivated by the fact that data distribution and tasks significantly vary across different recommendation scenarios. In the context of direct recommendation, we capitalize on the power of the instruction-tuned LLM to predict items that a user may like or dislike. We form pairs of these items to calculate the Bayesian Personalized Ranking (BPR) [46] loss. For sequential recommendation, we harness the capabilities of the instruction-tuned LLM to predict items that are highly preferred by the user. These predicted items are then randomly inserted into the sequence of items the user has interacted with. For rating prediction, we utilize the LLM to extract valuable side information (i.e., missing attributes), which is then seamlessly integrated as additional features within the training data.
III-B1 Data Augmentation for Direct Recommendation